OpenStudy (anonymous):

I'm not good at these, can someone please help me on.. 4(x-b)=x solve for x

jhonyy9 (jhonyy9):

so than 4x-4b=x than add to both sides (-x) and will get 4x-x-4b=x-x 3x-4b=0 add to both sides 4b 3x-4b+4b=4b 3x=4b divide both sides by 3 x=4b/3

OpenStudy (anonymous):

Oh i get it now, Thanks once again